# Lost in Benchmarks? Rethinking Large Language Model Benchmarking with Item Response Theory

### v2

@@ 1 Introduction
As the scale and performance of large language models (LLMs) continue to grow, accurately measuring their capabilities has become increasingly important. Currently, the performance of LLMs is primarily evaluated through various benchmarks, which are comprehensive test suites consisting of carefully designed questions to assess model behavior across different tasks. However, in practice, existing benchmarks often exhibit significant limitations, prompting consideration of their effectiveness (McIntosh et al. 2024).
@@
As illustrated in Figure 2, on one hand, different benchmarks often produce inconsistent leaderboards (Perlitz et al. 2024). This issue persists even for benchmarks designed to evaluate similar underlying capabilities, leading to the same model achieving substantially varying rankings across evaluations. On the other hand, many benchmarks show weak separability among top models, limiting the ability to understand performance differences and further model refinement. Given these limitations, there is a growing need for a more systematic analysis of LLM benchmarks.
In this paper, we conduct a comprehensive analysis for various popular LLM benchmark datasets. Our experiment is based on Item Response Theory (IRT) (Lord, Novick, and Birnbaum 1968; Baker and Kim 2004), a psychometric framework widely used in educational assessment, analyzes item properties such as difficulty by modeling item response against examinee ability [^3] In the paper, an item refers to a benchmark sample.. However, traditional IRT struggles with the complexity and scale of modern datasets, and its assumption of the normal distribution of abilities does not always hold.
@@
    We use PSN-IRT for an in-depth analysis of mainstream benchmarks, and found that current LLM benchmarks present deficiencies in many aspects.
@@
Figure 2: The illustration of our proposed PSN-IRT. Separate neural networks estimate model ability ($\theta$) and item parameters ($a,b,c,d$), which are then combined via the IRT formula to predict the probability of a correct response. After that, the networks can be leveraged for estimating properties for models or items, respectively.
@@ 2.1 Related Work
Item Response Theory (IRT) (Lord, Novick, and Birnbaum 1968; Baker and Kim 2004) offers a psychometric framework for data-driven, item-level analysis. Within NLP, IRT has been primarily utilized to diagnose fundamental item properties such as difficulty and discriminability (Vania et al. 2021; Byrd and Srivastava 2022). More recently, its principles have also been used to inspire methods for improving evaluation efficiency, for instance, through optimized item selection (Maia Polo et al. 2024) or adaptive testing paradigms (Zhuang et al. 2024). However, due to the complexity and scale of modern datasets and its assumption of normally distributed abilities, the application of IRT on LLM benchmarks is still underexplored (Ye et al. 2025).
@@ 2.2 Preliminary: Item Response Theory
Item Response Theory (IRT) (Lord, Novick, and Birnbaum 1968) is a psychometric framework widely used in educational and cognitive assessments to model the relationship between benchmark items and examinee abilities. Unlike Classical Test Theory (CTT) (DeVellis 2006), which relies on aggregate test scores, IRT analyzes individual item responses as a function of a latent ability $\theta$, enabling precise measurement of both item and examinee properties.
@@
More advanced IRT models introduce additional parameters to capture diverse item properties. However, traditional IRT often suffer from inaccurate parameter estimation and rely on assumptions like normal ability distributions, which may not align with real-world data (Tsutsumi, Kinoshita, and Ueno 2021).
@@ 3 Pseudo-Siamese Network for IRT
Upon completion of training, the two networks of PSN-IRT can be used for evaluating model performance and diagnosing benchmark characteristics, respectively. Specifically, the Model Network, by processing a model’s encoding, outputs an estimate of the model ability $\theta$. Meanwhile, the Item Network, given an benchmark item’s encoding, outputs its four estimated psychometric parameters $a,b,c,d$, which can be used for an in-depth diagnosis of individual item characteristics and the overall benchmark quality.
@@ 4.1 Experimental Setup
We mainly focus on currently high-performing models, including 360GPT2-Pro [^4] https://ai.360.cn/, DeepSeek-V3 (DeepSeek-AI et al. 2025), Doubao-Pro [^5] https://www.volcengine.com/product/doubao, Gemini-1.5 (Team et al. 2024a), Hunyuan-Turbo [^6] https://hunyuan.tencent.com/, Moonshot-v1777https://platform.moonshot.cn/, Qwen-Plus (Qwen et al. 2025), and Yi-Lightning (Wake et al. 2025). However, testing only strong models might obscure differences in the discriminability of the test sets, as most items could be solved easily, thus limiting a comprehensive evaluation of test set quality (Mart´ınez-Plumed et al. 2019). To address this, we intentionally included several relatively weaker models, including Gemma-2B-it (Team et al. 2024b), Mistral-7B-Instruct-v0.1 (Jiang et al. 2023), Qwen2.5-3B-Instruct (Qwen et al. 2025), and Vicuna-7B-v1.3 (Chiang et al. 2023).
@@
We mainly compare PSN-IRT with traditional IRT methods, namely IRT estimated based on traditional parameter estimation techniques, including Maximum Likelihood Estimation (MLE), Markov Chain Monte Carlo (MCMC) (Hastings 1970), Variational Inference (VI) (Jordan et al. 1999), and VIBO (Wu et al. 2020). We also compare with Deep-IRT (Tsutsumi, Kinoshita, and Ueno 2021), which is an extension of 1PL IRT that learns item-trait interactions via deep networks.
@@ 4.2 Main Results
Given our goal to comprehensively capture diverse item characteristics, we adopt the more expressive 4PL structure for all methods except Deep-IRT. Experiments also confirm that PSN-IRT performs best with 4PL, with detailed results provided in the Appendix B.1.
@@ 4.3 Ablation Study
To assess whether more complex architectures offer benefits over our straightforward design, we conduct an ablation study. Specifically, we compare PSN-IRT with two variants: one replaces one-hot inputs with semantic embeddings from Instructor (Su et al. 2023), and the other uses a GNN (Su et al. 2022) instead of the MLP backbone [^8] Please refer to Appendix A.2 for more details..
@@
Chinese SimpleQA 2 3 1 9 3 5 21
MBPP 5 6 4 8 4 4 26
MATH 4 1 2 7 7 10 27
TheoremQA 1 10 3 11 2 2 27
GPQA Diamond 3 11 6 10 1 1 30
MMLU 9 9 9 5 5 3 31
BBH 6 5 8 6 6 7 32
GSM8K 8 2 5 3 11 11 32
HellaSwag 10 7 10 2 8 6 33
HumanEval 7 4 7 4 9 9 33
ARC-C 11 8 11 1 10 8 38
Table 5: Ranks of datasets based on average item-level properties estimated by PSN-IRT. Each dataset’s score is the mean of its benchmark items’ property. Lower values indicate better performance in that property.
@@ 5.2 Benchmark Analysis
Based on the measurements, we provide an aggregate rank for each benchmark, calculated from the sum of individual ranks in Table 5, where a lower total score indicates better overall quality. Additionally, Figure 3 visualizes the distribution of item-level properties across the 11 LLM benchmarks. Building on these experimental results, our detailed analysis is presented as follows.
@@
As shown in Table 5, difficulty is varied among different benchmarks. Datasets such as TheoremQA, Chinese SimpleQA, and GPQA Diamond lead in difficulty and continue to challenge many LLMs. In contrast, other benchmarks including ARC-C, HellaSwag, and HumanEval have become comparatively easy, where many of the items are readily solved by high-performing models, diminishing their utility for differentiating top-tier systems.
Crucially, as shown in Figure 3, even within more challenging benchmarks, items often present only moderate difficulty to elite LLMs. For instance, item difficulty scores in the IRT scaling may be as low as 1.0 for models exhibiting ability scores above 3.0. This suggests that while such benchmarks can assess average models, they are less effective at separating most advanced systems, suggesting that they are insufficient for comprehensively evaluating frontier model capabilities.
@@
While Table 5 indicates that datasets like GPQA Diamond and TheoremQA achieve the highest relative ranks for average LEH, suggesting they offer more headroom than other benchmarks, the absolute LEH values in Figure 3 are often smaller than ideal. Conversely, benchmarks such as MATH and particularly GSM8K exhibit even lower average LEH scores. This signifies that current high-performing models are already approaching or have reached the performance ceiling for a substantial portion of items within these test sets. This necessitates the development of new benchmark items and datasets explicitly designed with greater headroom as models continue to evolve.
@@
The guessing-rate reflects the chance that a model can answer a question correctly without actual understanding, typically by exploiting format-based cues or shortcuts. As shown in Figure 3, datasets like ARC-C, HellaSwag, and MMLU generally present higher guessing-rates, possibly due to their multiple-choice formats. Notably, further inspection of the item guessing-rate parameter distributions, as depicted in Figure 3, reveals that a considerable number of items across several benchmarks exhibit unusually high guessing-rates. This observation regards the potential for such items to indicate data contamination, namely the content or answers for these items might have been present in the models’ pre-training data (Zhuang et al. 2024).
@@ 6 Efficient Benchmarking with PSN-IRT
To evaluate the effectiveness of different item selection strategies, we design an experiment to evaluate the effectiveness of different item selection strategies. Our methodology involve Benchmark Agreement Testing (BAT) (Perlitz et al. 2024). We first establish a reference model ranking by aggregating results from two public leaderboards: Chatbot Arena (Chiang et al. 2024) and OpenCompass Arena (Contributors 2023) [^9] These arenas are based on human evaluation and thus reflect human preference regarding overall model capabilities.. Subsequently, various subsets of benchmark items are chosen based on different criteria. Model rankings generated using these subsets are then compared against the reference arena ranking using Kendall’s $\tau$ to measure agreement. Alongside agreement, we also consider the variance of model scores on these subsets; higher variance generally indicates the subset’s capacity to differentiate more clearly between models. The evaluations are conducted using two distinct model groupings: one set comprising a mix of stronger and weaker models ("w/ weak models"), and another set from which weaker models are excluded ("w/o weak models"), to specifically assess separability among stronger contenders.
@@ Results.
In contrast, other item selection approaches are less effective for differentiating among strong models. For instance, selecting items based on discriminability, results in zero score variance when weak models are excluded, indicating it merely separates broad capability tiers rather than providing granularity among top-performers. Separately, a strategy of clustering items based on their success/failure vectors (Pacchiardi, Cheke, and Hernández-Orallo 2024) also fails to improve ranking correlation, even when it increases score separability.
@@ 7 Conclusion
In this work, we introduce PSN-IRT, a framework that combines psychometrics with neural networks to provide comprehensive and reliable analyses of LLM benchmarks. Through extensive experiments on 12 LLMs across 11 diverse datasets, we demonstrate that current benchmarks often suffer from uneven measurement properties, insufficient difficulty ceilings, item saturation, and data contamination. Moreover, we show that strategically selecting smaller, high-quality item sets using PSN-IRT enhances alignment with human preference and separability among top models.
@@ A.2 Ablation Study
where $\mathbf{e}_{q}$ is the item’s initial embedding, $\mathbf{W}_{rel}$ is a trainable weight matrix that projects the scalar response $rel$ into the embedding space, and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation. This combined vector is then processed by a single linear layer ($\text{Linear}_{\text{gate}}$) to form the cognitive degree. Subsequently, a cognitive state $\mathbf{c}_{s,q}$ is formed through an element-wise product:

### v3

@@ 1 Introduction
As the scale and performance of large language models (LLMs) continue to grow, accurately measuring their capabilities has become increasingly important (Chang et al. 2024; Zhou et al. 2024; Huang et al. 2025). Currently, the performance of LLMs is primarily evaluated through various benchmarks (Wang et al. 2024), which are comprehensive test suites consisting of carefully designed questions to assess model behavior across different tasks. However, in practice, existing benchmarks often exhibit significant limitations, prompting consideration of their effectiveness (McIntosh et al. 2024).
@@
As illustrated in Figure 2, on one hand, even benchmarks that are designed to measure similar underlying capabilities often produce inconsistent leaderboards, leading to substantial ranking variations for the same model (Perlitz et al. 2024). On the other hand, many benchmarks show weak separability among top models, limiting the ability to understand performance differences and further model refinement. Given these limitations, there is a growing need for a more systematic analysis of LLM benchmarks.
In this paper, we conduct a comprehensive analysis for various popular LLM benchmark datasets. Our experiment is based on Item Response Theory (IRT) (Lord et al. 1968; Baker and Kim 2004), a psychometric framework widely used in educational assessment to analyze item properties such as difficulty by modeling item response against examinee ability [^3] In the paper, an item refers to a benchmark sample.. However, traditional IRT struggles with the complexity and scale of modern datasets, and its assumption of the normal distribution of abilities does not always hold.
@@
    We use PSN-IRT for an in-depth analysis of mainstream benchmarks and find that current LLM benchmarks present deficiencies in many aspects.
@@
Figure 2: The illustration of our proposed PSN-IRT. Separate neural networks estimate model-ability ($\theta$) and item parameters ($a,b,c,d$), which are then combined via the IRT formula to predict the probability of a correct response. After that, the networks can be leveraged for estimating properties for models or items, respectively.
@@ 2.1 Related Work
Item Response Theory (IRT) (Lord et al. 1968; Baker and Kim 2004) offers a psychometric framework for data-driven, item-level analysis. Within natural language processing (NLP), IRT has been primarily utilized to diagnose fundamental item properties such as difficulty and discriminability (Vania et al. 2021; Byrd and Srivastava 2022). More recently, its principles have also been used to inspire methods for improving evaluation efficiency, for instance, through optimized item selection (Maia Polo et al. 2024) or adaptive testing paradigms (Zhuang et al. 2025; Truong et al. 2025). However, due to the complexity and scale of modern datasets and its assumption of normally distributed abilities, the application of IRT on LLM benchmarks is still underexplored (Ye et al. 2025).
@@ 2.2 Preliminary: Item Response Theory
Item Response Theory (IRT) (Lord et al. 1968) is a psychometric framework widely used in educational and cognitive assessments to model the relationship between benchmark items and examinee abilities. Unlike Classical Test Theory (CTT) (DeVellis 2006), which relies on aggregate test scores, IRT analyzes individual item responses as a function of a latent ability $\theta$, enabling precise measurement of both item and examinee properties.
@@
More advanced IRT models introduce additional parameters to capture diverse item properties. However, traditional IRT often suffer from inaccurate parameter estimation and rely on assumptions like normal ability distributions, which may not align with real-world data (Tsutsumi et al. 2021).
@@ 3 Pseudo-Siamese Network for IRT
Notably, PSN-IRT follows standard IRT assumptions such as unidimensionality and monotonicity, which generally hold for LLM benchmarking (Kipnis et al. 2025; Truong et al. 2025).
@@
Upon completion of training, the two networks of PSN-IRT can be used for evaluating model performance and diagnosing benchmark characteristics, respectively. Specifically, the model network, by processing a model’s encoding, outputs an estimate of the model-ability $\theta$. Meanwhile, the item network, given a benchmark item’s encoding, outputs its four estimated psychometric parameters $a,b,c,d$, which can be used for an in-depth diagnosis of individual item characteristics and the overall benchmark quality.
@@ 4.1 Experimental Setup
We mainly focus on currently high-performing models, including 360GPT2-Pro [^4] https://ai.360.cn/, DeepSeek-V3555https://platform.deepseek.com/, Doubao-pro [^6] https://www.volcengine.com/product/doubao, Gemini-1.5777https://ai.google.dev/, Hunyuan-Turbo [^8] https://hunyuan.tencent.com/, Moonshot-v1999https://platform.moonshot.cn/, Qwen-Plus [^10] https://qwen.ai/apiplatform, and Yi-Lightning [^11] https://platform.lingyiwanwu.com/. However, testing only strong models might obscure differences in the discriminability of the test sets, as most items could be solved easily, thus limiting a comprehensive evaluation of test set quality (Martínez-Plumed et al. 2019). To address this, we intentionally included several relatively weaker models, including Gemma-2B-it (Team et al. 2024), Mistral-7B-Instruct-v0.1 (Jiang et al. 2023), Qwen2.5-3B-Instruct (Qwen et al. 2025), and Vicuna-7B-v1.3 (Chiang et al. 2023).
@@
We mainly compare PSN-IRT with traditional IRT methods, namely IRT estimated based on traditional parameter estimation techniques, including Maximum Likelihood Estimation (MLE), Markov Chain Monte Carlo (MCMC) (Hastings 1970), Variational Inference (VI) (Jordan et al. 1999), and VIBO (Wu et al. 2020). We also compare with Deep-IRT (Tsutsumi et al. 2021), which is an extension of 1PL IRT that learns item-trait interactions via deep networks.
@@ 4.2 Main Results
Given our goal to comprehensively capture diverse item characteristics, we adopt the more expressive 4PL structure for all methods except Deep-IRT. Experiments also confirm that PSN-IRT performs best with 4PL.
@@ 4.3 Ablation Study
To assess whether more complex architectures offer benefits over our straightforward design, we conduct an ablation study. Specifically, we compare PSN-IRT with two variants: one replaces one-hot inputs with semantic embeddings from Instructor (Su et al. 2023), and the other uses a GNN (Su et al. 2022) instead of the MLP backbone.
@@
Chinese SimpleQA 2 3 1 9 3 5 23
TheoremQA 1 10 3 11 2 2 29
MBPP 5 6 4 8 4 4 31
MATH 4 1 2 7 7 10 31
GPQA Diamond 3 11 6 10 1 1 32
BBH 6 5 8 6 6 7 38
MMLU 9 9 9 5 5 3 40
GSM8K 8 2 5 3 11 11 40
HumanEval 7 4 7 4 9 9 40
HellaSwag 10 7 10 2 8 6 43
ARC-C 11 8 11 1 10 8 49
Table 5: Ranks of datasets based on average item-level properties estimated by PSN-IRT. Each rank is obtained by averaging the property over all items in a dataset and then sorting these averages; smaller rank values correspond to better performance.
@@ 5.2 Benchmark Analysis
Based on the measurements, we provide an aggregate rank for each benchmark, calculated from the sum of individual ranks in Table 5, where a lower total rank value indicates better overall quality. Additionally, Figure 3 visualizes the distribution of item-level properties across the 11 LLM benchmarks. Building on these experimental results, our detailed analysis is presented as follows.
@@
As shown in Table 5, difficulty is varied among different benchmarks. Datasets such as TheoremQA, Chinese SimpleQA, and GPQA Diamond lead in difficulty and continue to challenge many LLMs. In contrast, other benchmarks including ARC-C, HellaSwag, and MMLU have become comparatively easy, where many of the items are readily solved by high-performing models, diminishing their utility for differentiating top-tier systems.
Crucially, as shown in Figure 3, even the most challenging items in existing benchmarks exhibit relatively low difficulty levels. For instance, on the IRT scale, the highest item difficulty values rarely exceed 1.0, while top models such as DeepSeek-V3 have estimated abilities well above 3.0. This indicates that even the hardest questions are significantly below the capability level of elite LLMs, highlighting a lack of sufficient challenge for frontier capabilities.
@@
While Table 5 indicates that datasets like GPQA Diamond and TheoremQA achieve the highest relative ranks for average LEH, suggesting they offer more headroom than other benchmarks, the absolute LEH values in Figure 3 are often smaller than ideal. Conversely, benchmarks such as GSM8K and ARC-C exhibit lower average LEH scores. This signifies that current high-performing models are already approaching or have reached the performance ceiling for a substantial portion of items within these test sets. This necessitates the development of new benchmark items and datasets explicitly designed with greater headroom as models continue to evolve.
@@
The guessing-rate reflects the chance that a model can answer a question correctly without actual understanding, typically by exploiting format-based cues or shortcuts. As shown in Figure 3, datasets like ARC-C, HellaSwag, and MMLU generally present higher guessing-rates, possibly due to their multiple-choice formats. Notably, further inspection of the item guessing-rate parameter distributions, as depicted in Figure 3, reveals that a considerable number of items across several benchmarks exhibit unusually high guessing-rates. This observation regards the potential for such items to indicate data contamination, namely the content or answers for these items might have been present in the models’ pre-training data (Zhuang et al. 2025).
@@ 6 Efficient Benchmarking with PSN-IRT
To evaluate the effectiveness of different item selection strategies, we design an experiment to evaluate the effectiveness of different item selection strategies. Our methodology involves Benchmark Agreement Testing (BAT) (Perlitz et al. 2024). We first establish a reference model ranking by aggregating results from two public leaderboards: Chatbot Arena (Chiang et al. 2024) and OpenCompass Arena (Contributors 2023) [^12] These arenas are based on human evaluation and thus reflect human preference regarding overall model capabilities..
Subsequently, various subsets of benchmark items are chosen based on different criteria. Model rankings generated using these subsets are then compared against the reference arena ranking using Kendall’s $\tau$ to measure agreement. Alongside agreement, we also consider the variance of model scores on these subsets; higher variance generally indicates the subset’s capacity to differentiate more clearly between models. The evaluations are conducted using two distinct model groupings: one set comprising a mix of stronger and weaker models ("w/ weak models"), and another set from which weaker models are excluded ("w/o weak models"), to specifically assess separability among stronger contenders, with weak models introduced in Section 4.1.
@@ Results.
In contrast, other item selection approaches are less effective for distinguishing strong models. For instance, selecting items based on discriminability yields zero score variance when weak models are excluded, indicating it merely separates broad capability tiers rather than providing granularity among top-performers. Separately, clustering items based on their success/failure vectors (Pacchiardi et al. 2024) also fails to improve ranking correlation, even when it increases score separability.
@@ 7 Conclusion
In this work, we introduce PSN‑IRT, a framework that combines psychometrics with neural networks to provide comprehensive and reliable analyses of LLM benchmarks. Through extensive experiments on 12 LLMs across 11 diverse datasets, we demonstrate that current benchmarks often suffer from uneven measurement properties, insufficient difficulty ceilings, item saturation, and data contamination. Moreover, we show that strategically selecting smaller, high‑quality item sets using PSN‑IRT enhances alignment with human preference and separability among top models.
Acknowledgments
This work is supported by National Natural Science Foundation of China (62276077, 62376075, 62376076), Department of Science and Technology of Heilongjiang (Grant No. ZY04JD04), and the Key Laboratory of Computing Power Network and Information Security, Ministry of Education (Grant No. 2023ZD027).
@@ A.2 Ablation Study
where $\mathbf{e}_{q}$ is the item’s initial embedding, $\mathbf{W}_{rel}$ is a trainable weight matrix that projects the scalar response relrel into the embedding space, and $\cdot$$\cdot$] denotes concatenation. This combined vector is then processed by a single linear layer ($\text{Linear}_{\text{gate}}$) to form the cognitive degree. Subsequently, a cognitive state $\mathbf{c}_{s,q}$ is formed through an element-wise product: